two hikers, Alan and Ben, are 22.2 miles apart and walking towards each other. They meet in 6 hours. Find the rate of each hiker if ben walks 0.9 mph faster than Alan
Answer by TimothyLamb(4379)

x = speed of alan (mph)
y = speed of ben (mph)
---
y = x + 0.9
---
speed of hikers relative to each other:
s = d/t
s = 22.2/6
s = x + y
22.2/6 = x + y
---
linear system:
y = x + 0.9
22.2/6 = x + y
---
put the system of linear equations into standard form
---
x - y = -0.9
x + y = 3.7
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1.4
y = 2.3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---